# Entanglement measures for detectability

### v2

@@ IV Information-geometrical characterization
Then, using the operator $(\log\sigma_{AB}^{*}-\log\rho_{AB})$, we define the hyperplane ${\cal M}(\sigma_{AB}^{*},\rho_{AB})$ as (See Fig. 1.)
@@
Figure 1: Entanglement witness: This figure shows the relation among the entangled state $\rho_{AB}$, the set of separable states ${\cal S}_{AB}$, and the separation hyperplane ${\cal M}(\sigma_{AB}^{*},\rho_{AB})$.
The following theorem guarantees that ${\cal M}(\sigma_{AB}^{*},\rho_{AB})$ separate $\rho_{AB}$ from ${\cal S}_{AB}$. Hence, the operator $(\log\sigma_{AB}^{*}-\log\rho_{AB})-D(\sigma_{AB}^{*}\|\rho_{AB})$ works as entanglement witness for $\rho_{AB}$.
@@
(i) For any separable state $\sigma_{AB}\in{\cal S}$, we have
@@
To grasp the geometrical meaning of the above fact, we employ information geometry. Given a Hermitian matrix $X$, we often focus on the set :=${\cal E}:=\{\rho_{t}\}$, where $\rho_{t}:=\exp(tX+\log\rho_{0})/c_{t}$ with $c_{t}:=\operatorname{\mathrm{Tr}}\exp(tX+\log\rho_{0})$. The set $\{\rho_{t}\}$ is called an exponential family generated by $X$. Also, the set :=${\cal M}:=\{\rho|\operatorname{\mathrm{Tr}}\rho X=\operatorname{\mathrm{Tr}}\rho_{0}X\}$ is called a mixture family generated by $X$. Any elements $\sigma\in{\cal M}$ and $\rho\in{\cal E}$ satisfy
$D(\sigma\|\rho)=D(\sigma\|\rho_{0})+D(\rho_{0}\|\rho),$ (23)
which is called Pythagorean theorem [13], [42, Theorem 2.3]. We apply (23) to the case with $X=-(\log\sigma_{AB}^{*}-\log\rho_{AB})$ and $\rho_{0}=\sigma_{AB}^{*}$. Any element $\sigma_{AB}^{\prime}\in{\cal M}(\sigma_{AB}^{*},\rho_{AB})$ satisfies
@@
Then, ${\cal M}(\sigma_{AB}^{*},\rho_{AB})$ is the tangent space of the sphere ${\cal S}_{D}$ at $\sigma_{AB}^{*}$. If a separable state $\sigma_{AB}$ satisfies
@@ V Coherence measure
For further analysis on our entanglement measures, we study the coherence measures defined below, because, as explained in the next section, our entanglement measures are related to these coherence measures. We consider a state $\sum_{i,j}\theta_{i,j}|j\rangle\langle i|$ on a $d$-dimensional Hilbert ${\mathcal{H}}$ spanned by $\{|j\rangle\}_{j=0}^{d-1}$, and denote the set of incoherent states by ${\cal I}_{c}$, which is written as $\{\sigma(p)\}$ with $\sigma(p):=\sum_{i}p_{i}|i\rangle\langle i|$, where $p=(p_{i})$ is a distribution. Then, we define the following coherence measures.
@@
In the following discussion, we employ the operators $Z:=\sum_{j=0}^{d-1}\omega^{j}|j\rangle\langle j|$ with $\omega:=e^{2\pi\sqrt{-1}/d}$ and $X:=\sum_{j=0}^{d-1}|j+1\rangle\langle j|$. We define $\rho(p):=\sum_{j=0}^{d-1}p_{j}Z^{j}|+\rangle\langle+|Z^{-j}$ with $|+\rangle:=\frac{1}{\sqrt{d}}\sum_{j=0}^{d-1}|j\rangle$. Then, we have
@@
where $p_{mix}$ is the uniform distribution. Since the above second equations follow from the direct calculation, the following analysis shows the first equations.
@@
where $\rho_{mix}$ is the completely mixed state. The above relations imply the first equation (31). Similarly, since $x\mapsto x^{\alpha}$ is matrix convex for $\alpha>1$,
@@
which implies the first equation (32) for $\alpha>1$. Similarly, since $x\mapsto x^{\alpha}$ is matrix concave for $\alpha\in[0,1)$, an analysis similar to (34) yields the first equation (32) for $\alpha\in[0,1)$.
@@
$=$ $\left\{\begin{array}[]{l}\max_{\tau}\frac{\alpha}{\alpha-1}\log\|(\rho^{\frac{1-\alpha}{2\alpha}}\tau^{-\frac{1-\alpha}{\alpha}}\rho^{\frac{1-\alpha}{2\alpha}})^{\rm diag}\|\\ \hskip 150.69397pt\hbox{ for }\alpha\in(0,1)\\ \max_{\tau}\frac{-\alpha}{\alpha-1}\log\|((\rho^{\frac{1-\alpha}{2\alpha}}\tau^{-\frac{1-\alpha}{\alpha}}\rho^{\frac{1-\alpha}{2\alpha}})^{\rm diag})^{-1}\|\\ \hskip 150.69397pt\hbox{ for }\alpha>1\end{array}\right.$ (41)
@@
Proof: The idea of this proof comes from the proof of [48, Lemma 6 and (B4)]. We choose $\beta$ such that $\frac{1}{\beta}+\frac{1}{\alpha}=1$ so that $\beta=\frac{\alpha-1}{\alpha}$. We fix $\alpha\in(0,1)$ so that $\beta<0$. For any state $\tau$ and any incoherent state $\sigma\in{\cal I}_{c}$, the reverse matrix Hölder inequality [51, Theorem 2.7] implies
@@
Here, $(a)$ follows from the minimax theorem because $\operatorname{\mathrm{Tr}}\rho^{\frac{1-\alpha}{2\alpha}}\tau^{\beta}\rho^{\frac{1-\alpha}{2\alpha}}\sigma$ is linear for $\sigma$ and is convex for $\tau$. Hence, we obtain the desired equation with $\alpha\in(0,1)$.
Next, we fix $\alpha>1$ so that $01>\beta>$ 0. The matrix Hölder inequality [52],[51, Theorem 2.6] implies
@@
hold.
Proof: Since the first equation in (48) follows from the second equation in (48) with taking the limit $\alpha\to 1$, we show only the second equation in (48). The idea of this proof is similar to the proof of [48, Lemma 7]. For $\alpha\in(0,1)$, using (41), we have
@@
$=$ $\max_{\tau}\frac{\alpha}{\alpha-1}\log\|((\rho_{1}\otimes\rho_{2})^{\frac{1-\alpha}{2\alpha}}\tau^{-\frac{1-\alpha}{\alpha}}(\rho_{1}\otimes\rho_{2})^{\frac{1-\alpha}{2\alpha}})^{\rm diag}\|$
$\geq$ $\max_{\tau_{1},\tau_{2}}\frac{\alpha}{\alpha-1}\log\|((\rho_{1}\otimes\rho_{2})^{\frac{1-\alpha}{2\alpha}}(\tau_{1}\otimes\tau_{2})^{-\frac{1-\alpha}{\alpha}}$
$\hskip 86.11084pt\cdot(\rho_{1}\otimes\rho_{2})^{\frac{1-\alpha}{2\alpha}})^{\rm diag}\|$
$=$ $\max_{\tau_{1},\tau_{2}}\frac{\alpha}{\alpha-1}(\log\|(\rho_{1}^{\frac{1-\alpha}{2\alpha}}\tau_{1}^{-\frac{1-\alpha}{\alpha}}\rho_{1}^{\frac{1-\alpha}{2\alpha}})^{\rm diag}\|$
$+\log\|(\rho_{2}^{\frac{1-\alpha}{2\alpha}}\tau_{2}^{-\frac{1-\alpha}{\alpha}}\rho_{2}^{\frac{1-\alpha}{2\alpha}})^{\rm diag}\|)$
@@
Table 2: List of mathematical symbols for coherence.
@@
$I_{c}(\rho)$ Closet incoherent state (35)
@@ VI Maximally correlated states
The discussion in Section II shows the operational meaning of $E_{\alpha}^{*}(\rho_{AB})$ including $E^{*}(\rho_{AB})$. We calculate $E^{*}(\rho_{AB})$ and $E^{*}_{\alpha}(\rho_{AB})$ for several special cases. When $\rho_{AB}$ is a pure state $|\phi\rangle\langle\phi|$ $E^{*}(\rho_{AB})$ and $E^{*}_{\alpha}(\rho_{AB})$ with $\alpha>1$ are infinity. However, $E^{*}_{\alpha}(\rho_{AB})$ with $\alpha\in[0,1)$ takes a finite value as follows. When the Schmidt coefficients of $|\phi\rangle\langle\phi|$ are $\sqrt{p_{i}}$ with $i=1,\ldots,d$, $\max_{\sigma_{AB}\in{\cal S}_{AB}}\langle\phi|\sigma|\phi\rangle=\max_{i}p_{i}$ [44, (40) and (97)]. Hence,
$e^{(\alpha-1)E_{\alpha}^{*}(\rho_{AB})}=\max_{\sigma_{AB}\in{\cal S}_{AB}}\operatorname{\mathrm{Tr}}(\rho^{\frac{1-\alpha}{2\alpha}}\sigma_{AB}\rho^{\frac{1-\alpha}{2\alpha}})^{\alpha}$
$=$ $\max_{\sigma_{AB}\in{\cal S}_{AB}}\operatorname{\mathrm{Tr}}(|\phi\rangle\langle\phi|\sigma_{AB}|\phi\rangle\langle\phi|)^{\alpha}=\max_{\sigma_{AB}\in{\cal S}_{AB}}\langle\phi|\sigma_{AB}|\phi\rangle^{\alpha}$
$=$ $(\max_{\sigma_{AB}\in{\cal S}_{AB}}\langle\phi|\sigma_{AB}|\phi\rangle)^{\alpha}=(\max_{i}p_{i})^{\alpha}.$ (50)
@@
More generally, the calculation of these quantities can be converted to the calculation of the coherence measure for special states $\rho_{AB}$. Similar to the paper [44], we consider maximally correlated states. A state $\rho_{AB}$ on ${\mathcal{H}}_{A}\otimes{\mathcal{H}}_{B}$ is called a maximally correlated state when there exist bases $\{|u_{j,A}\rangle\}_{j}$ and $\{|u_{j,B}\rangle\}_{j}$ of ${\mathcal{H}}_{A}$ and ${\mathcal{H}}_{B}$ such that the support $\rho_{AB}$ is included in the subspace spanned by $\{|u_{j,A}\rangle|u_{j,B}\rangle\}_{j}$ on ${\mathcal{H}}_{A}\otimes{\mathcal{H}}_{B}$ [49, 50], [42, Section 8.8]. In this case, the state $\rho_{AB}$ is written as $\sum_{i,j}\theta_{i,j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{i,A}|\langle u_{i,B}|$. In this case, the calculation of $E^{*}(\rho_{AB})$ and $E^{*}_{\alpha}(\rho_{AB})$ can be simplified as follows. Similar to Theorem 1 in [44], the calculation of $E^{*}(\rho_{AB})$ and $E^{*}_{\alpha}(\rho_{AB})$ are reduced to coherence measures.
@@
where $\rho_{\theta}:=\sum_{i,j}\theta_{i,j}|j\rangle\langle i|$. The combination of Lemmas 6 and 7 guarantees that any maximally entangled state $\rho_{AB}$ satisfies
@@
$=\frac{1}{\alpha-1}\log\sum_{j=0}^{d-1}p_{j}^{1-\alpha}-\frac{\alpha}{\alpha-1}\log d.$ (57)
However, it is not easy to calculate these quantities for general states. Therefore, it is needed to develop an efficient method for their calculation in the general case.

### v3

@@ IV-A Single system
That is, the minimizer of $E^{*}(\rho_{AB})$ produces entanglement witness. In addition, we propose an algorithm to calculate our entanglement measure $E^{*}(\rho_{AB})$ in Section VII. As explained in Section VIII-B, the combination between the algorithm for $\sigma_{AB}^{*}$ and a procedure to find an entanglement witness based on quantum relative entropy enables us to derive an algorithm to find an entanglement witness. Hence, the relation between entanglement witness and the minimizer of $E^{*}(\rho_{AB})$ is quite important.
Although quantum relative entropy is a two-input quantity, it does not form an Euclidean geometry because it is not symmetric with respect to exchanging two input states. Hence, when we employ our entanglement measure, we need to employ a geometric structure different from the Euclidean geometry.
Using the operator $(\log\sigma_{AB}^{*}-\log\rho_{AB})$, we define the hyperplane ${\cal M}(\sigma_{AB}^{*},\rho_{AB})$ as (See Fig. 1.)
@@
Figure 1: Entanglement witness: This figure shows the relation among the entangled state $\rho_{AB}$, the set of separable states ${\cal S}_{AB}$, and the separation hyperplane ${\cal M}(\sigma_{AB}^{*},\rho_{AB})$. Any other element $\tilde{\sigma}_{AB}$ on the tangent space has larger relative entropy to $\rho_{AB}$ than the projected element ${\sigma}_{AB}^{*}$ due to the Pythagorean theorem (25).
The following theorem guarantees that ${\cal M}(\sigma_{AB}^{*},\rho_{AB})$ separates $\rho_{AB}$ from ${\cal S}_{AB}$. Hence, the operator $(\log\sigma_{AB}^{*}-\log\rho_{AB})-D(\sigma_{AB}^{*}\|\rho_{AB})$ works as an entanglement witness for $\rho_{AB}$. In fact, as discussed in [61, Eq. (4)], any entanglement witness can be constructed by a combination of local measurement because any matrix on the composite system ${\cal H}_{A}\otimes{\cal H}_{B}$ can be written as the combination of at most $(\dim{\cal H}_{A})^{2}(\dim{\cal H}_{B})^{2}$ matrices with a product form.
@@
(i) For any separable state $\sigma_{AB}\in{\cal S}_{AB}$, we have
@@
To grasp the geometrical meaning of the above fact, we employ information geometry. Given a Hermitian matrix $X$ and a state $\tilde{\rho}_{0}$, we often focus on the set :=${\cal E}:=\{\tilde{\rho}_{t}\}$, where $\tilde{\rho}_{t}:=\exp(tX+\log\tilde{\rho}_{0})/c_{t}$ with $c_{t}:=\operatorname{\mathrm{Tr}}\exp(tX+\log\tilde{\rho}_{0})$. The set =${\cal E}=\{\tilde{\rho}_{t}\}$ is called an exponential family generated by $X$. Also, the set :=${\cal M}:=\{\rho|\operatorname{\mathrm{Tr}}\rho X=\operatorname{\mathrm{Tr}}\tilde{\rho}_{0}X\}$ is called a mixture family generated by $X$.
Since the intersection ${\cal M}\cap{\cal E}$ is composed of only one element $\tilde{\rho}_{0}$, any elements $\sigma\in{\cal M}$ and $\rho\in{\cal E}$ satisfy
$D(\sigma\|\rho)=D(\sigma\|\tilde{\rho}_{0})+D(\tilde{\rho}_{0}\|\rho),$ (25)
which is called Pythagorean theorem [23], [12, Theorem 2.3].
We apply (25) to the case with $X=-(\log\sigma_{AB}^{*}-\log\rho_{AB})$ and $\tilde{\rho}_{0}=\sigma_{AB}^{*}$. Any element $\sigma_{AB}^{\prime}\in{\cal M}(\sigma_{AB}^{*},\rho_{AB})$ satisfies
@@
Since any state $\sigma_{AB}^{\prime}$ in ${\cal M}(\sigma_{AB}^{*},\rho_{AB})\setminus\{\sigma_{AB}^{*}\}$ satisfies $D(\sigma_{AB}^{\prime}\|\rho_{AB})=D(\sigma_{AB}^{\prime}\|\sigma_{AB}^{*})+D(\sigma_{AB}^{*}\|\rho_{AB})>D(\sigma_{AB}^{*}\|\rho_{AB})$ due to (25), any state $\sigma_{AB}^{\prime\prime}$ in the sphere ${\cal S}_{D}$ satisfies the condition $\operatorname{\mathrm{Tr}}\sigma_{AB}^{\prime\prime}(\log\sigma_{AB}^{*}-\log\rho_{AB})\leq D(\sigma_{AB}^{*}\|\rho_{AB})$. Hence, the hyperplane ${\cal M}(\sigma_{AB}^{*},\rho_{AB})$ is the tangent space of the sphere ${\cal S}_{D}$ at $\sigma_{AB}^{*}$.
If a separable state $\sigma_{AB}$ satisfies
@@
IV-B Simple proof of additivity
Next, we consider the composite system of two systems ${\cal H}_{A,1}\otimes{\cal H}_{B,1}$ and ${\cal H}_{A,2}\otimes{\cal H}_{B,2}$. We consider the states $\rho_{1}$ and $\rho_{2}$ on the respective systems. The following proposition is known.
Propositon 4 ([55, Proposition 2])
$E^{*}(\rho_{1}\otimes\rho_{2})=E^{*}(\rho_{1})+E^{*}(\rho_{2}).$ (30)
This proposition can be easily shown by using information geometry as follows.
We denote the sets of separable states on ${\cal H}_{A,j}\otimes{\cal H}_{B,j}$ and $({\cal H}_{A,1}\otimes{\cal H}_{A,2})\otimes({\cal H}_{B,1}\otimes{\cal H}_{B,2})$ by ${\cal S}_{AB,j}$ and ${\cal S}_{AB,12}$, respectively, for $j=1,2$. We choose $\sigma_{j}^{*}:=\mathop{\rm argmin}\limits_{\sigma_{j}\in{\cal S}_{AB,j}}D(\sigma_{j}\|\rho_{j})$. Then, $X_{j}:=(\log\sigma_{j}^{*}-\log\rho_{j})-D(\sigma_{j}^{*}\|\rho_{j})$ is an entanglement witness for $\rho_{j}$. That is, due to Theorem 3, any separable state $\sigma_{j}\in{\cal S}_{AB,j}$ satisfies
$\operatorname{\mathrm{Tr}}\sigma_{j}X_{j}\geq 0.$ (31)
Lemma 5
$X_{1}\otimes I+I\otimes X_{2}$ is an entanglement witness for $\rho_{1}\otimes\rho_{2}$.
Proof: For any element $\sigma\in{\cal S}_{AB,1,2}$, we have
$\operatorname{\mathrm{Tr}}\sigma(X_{1}\otimes I+I\otimes X_{2})$
$=$ $\operatorname{\mathrm{Tr}}(\operatorname{\mathrm{Tr}}_{2}\sigma)X_{1}+(\operatorname{\mathrm{Tr}}_{1}\sigma)X_{2}\geq 0,$ (32)
where the inequality follows from (31).
Proof of Proposition 4:   Now, we consider the mixture family ${\cal M}_{12}:=\{\rho|\operatorname{\mathrm{Tr}}\rho(X_{1}\otimes I+I\otimes X_{2})=0\}$ generated by the matrix $(X_{1}\otimes I+I\otimes X_{2})$. We often focus on the exponential family ${\cal E}_{12}:=\{\rho_{12,t}\}$ generated by $X$, where $\rho_{12,t}:=\exp(t(X_{1}\otimes I+I\otimes X_{2})+\log(\rho_{1}\otimes\rho_{2}))/c_{t}$ with $c_{12,t}:=\operatorname{\mathrm{Tr}}\exp(t(X_{1}\otimes I+I\otimes X_{2})+\log(\rho_{1}\otimes\rho_{2}))$.
Since the intersection ${\cal M}_{12}\cap{\cal E}_{12}$ is composed of only one element $\rho_{12,1}=\sigma_{1}^{*}\otimes\sigma_{2}^{*}$, Eq. (25) (Pythagorean theorem [23], [12, Theorem 2.3]) guarantees that any element $\sigma\in{\cal M}_{12}$ satisfies
$D(\sigma\|\rho_{1}\otimes\rho_{2})=$ $D(\sigma\|\sigma_{1}^{*}\otimes\sigma_{2}^{*})+D(\sigma_{1}^{*}\otimes\sigma_{2}^{*}\|\rho_{1}\otimes\rho_{2})$
$\geq$ $D(\sigma_{1}^{*}\otimes\sigma_{2}^{*}\|\rho_{1}\otimes\rho_{2}).$ (33)
Since any separable state $\sigma^{\prime}\in{\cal S}_{AB,12}$ is located in the opposite side of $\rho_{1}\otimes\rho_{2}$ with respect to hyperplane ${\cal M}_{12}$, the relation (33) yields
$D(\sigma^{\prime}\|\rho_{1}\otimes\rho_{2})\geq D(\sigma_{1}^{*}\otimes\sigma_{2}^{*}\|\rho_{1}\otimes\rho_{2}),$ (34)
which implies
$\min_{\sigma^{\prime}\in{\cal S}_{AB,12}}D(\sigma^{\prime}\|\rho_{1}\otimes\rho_{2})=D(\sigma_{1}^{*}\otimes\sigma_{2}^{*}\|\rho_{1}\otimes\rho_{2}).$ (35)
Therefore, we obtain (30). ∎
@@ V Coherence measure
Next, we consider how to calculate the proposed measure for maximally correlated states, which are special cases. For its preparation, we study the coherence measures defined below, because, as explained in the next section, our entanglement measures are related to these coherence measures. In particular, the analysis of maximally correlated states can be reduced to the analysis of the coherence measures. We consider a state $\sum_{i,j}\theta_{i,j}|j\rangle\langle i|$ on a $d$-dimensional Hilbert space ${\mathcal{H}}$ spanned by $\{|j\rangle\}_{j=0}^{d-1}$, and denote the set of incoherent states by ${\cal I}_{c}$, which is written as $\{\sigma(p)\}$ with $\sigma(p):=\sum_{i}p_{i}|i\rangle\langle i|$, where $p=(p_{i})$ is a distribution. Then, we define the following coherence measures.
@@
In the following discussion, we employ the operators $Z:=\sum_{j=0}^{d-1}\omega^{j}|j\rangle\langle j|$ with $\omega:=e^{2\pi\sqrt{-1}/d}$ and $X:=\sum_{j=0}^{d-1}|j+1\rangle\langle j|$. We define $\rho(p):=\sum_{j=0}^{d-1}p_{j}Z^{j}|+\rangle\langle+|Z^{-j}$ with $|+\rangle:=\frac{1}{\sqrt{d}}\sum_{j=0}^{d-1}|j\rangle$. Then, we have the following lemma.
Lemma 6
Using the uniform distribution $p_{mix}$, we have
@@
for $\alpha\in(0,1)\cup(1,2]$.
Proof: Since the above second equations follow from the direct calculation, the following analysis shows the first equations.
@@
where $\rho_{mix}$ is the completely mixed state. The above relations imply the first equation (39). Similarly, since $x\mapsto x^{\alpha}$ is matrix convex for $\alpha\in[1,2]$,
@@
which implies the first equation (40) for $\alpha\in(1,2]$. Similarly, since $x\mapsto x^{\alpha}$ is matrix concave for $\alpha\in[0,1)$, an analysis similar to (42) yields the first equation (40) for $\alpha\in[0,1)$.
@@
$=$ $\left\{\begin{array}[]{l}\max_{\tau}\frac{\alpha}{\alpha-1}\log\|(\rho^{\frac{1-\alpha}{2\alpha}}\tau^{-\frac{1-\alpha}{\alpha}}\rho^{\frac{1-\alpha}{2\alpha}})^{\rm diag}\|\\ \hskip 150.69397pt\hbox{ for }\alpha\in[1/2,1)\\ \max_{\tau}\frac{-\alpha}{\alpha-1}\log\|((\rho^{\frac{1-\alpha}{2\alpha}}\tau^{-\frac{1-\alpha}{\alpha}}\rho^{\frac{1-\alpha}{2\alpha}})^{\rm diag})^{-1}\|\\ \hskip 150.69397pt\hbox{ for }\alpha>1\end{array}\right.$ (49)
@@
Proof: The idea of this proof comes from the proof of [67, Lemma 6 and (B4)]. We choose $\beta$ such that $\frac{1}{\beta}+\frac{1}{\alpha}=1$ so that $\beta=\frac{\alpha-1}{\alpha}$. We fix $\alpha\in[1/2,1)$ so that $\beta\in[-1,0)$. For any state $\tau$ and any incoherent state $\sigma\in{\cal I}_{c}$, the reverse matrix Hölder inequality [70, Theorem 2.7] implies
@@
Here, $(a)$ follows from the minimax theorem because $\operatorname{\mathrm{Tr}}\rho^{\frac{1-\alpha}{2\alpha}}\tau^{\beta}\rho^{\frac{1-\alpha}{2\alpha}}\sigma$ is linear for $\sigma$ and is convex for $\tau$, which follows from the matrix convexity of $\tau\mapsto\tau^{\beta}$ with $\beta\in[-1,0)$. Hence, we obtain the desired equation with $\alpha\in(0,1)$.
Next, we fix $\alpha>1$ so that $01>\beta>$ 0. The matrix Hölder inequality [71],[70, Theorem 2.6] implies
@@
hold with $\alpha\in[1/2,1)\cup(1,\infty)$.
Proof: Since the first equation in (56) follows from the second equation in (56) with taking the limit $\alpha\to 1$, we show only the second equation in (56). The idea of this proof is similar to the proof of [67, Lemma 7]. For $\alpha\in(0,1)$, using (49), we have
@@
$=$ $\max_{\tau}\frac{\alpha}{\alpha-1}\log\big{\|}((\rho_{1}\otimes\rho_{2})^{\frac{1-\alpha}{2\alpha}}\tau^{-\frac{1-\alpha}{\alpha}}(\rho_{1}\otimes\rho_{2})^{\frac{1-\alpha}{2\alpha}})^{\rm diag}\big{\|}$
$\geq$ $\max_{\tau_{1},\tau_{2}}\frac{\alpha}{\alpha-1}\log\big{\|}((\rho_{1}\otimes\rho_{2})^{\frac{1-\alpha}{2\alpha}}(\tau_{1}\otimes\tau_{2})^{-\frac{1-\alpha}{\alpha}}$
$\hskip 86.11084pt\cdot(\rho_{1}\otimes\rho_{2})^{\frac{1-\alpha}{2\alpha}})^{\rm diag}\big{\|}$
$=$ $\max_{\tau_{1},\tau_{2}}\frac{\alpha}{\alpha-1}\big{(}\log\big{\|}(\rho_{1}^{\frac{1-\alpha}{2\alpha}}\tau_{1}^{-\frac{1-\alpha}{\alpha}}\rho_{1}^{\frac{1-\alpha}{2\alpha}})^{\rm diag}\big{\|}$
$+\log\big{\|}(\rho_{2}^{\frac{1-\alpha}{2\alpha}}\tau_{2}^{-\frac{1-\alpha}{\alpha}}\rho_{2}^{\frac{1-\alpha}{2\alpha}})^{\rm diag}\big{\|}\big{)}$
@@
TABLE II: List of mathematical symbols for coherence.
@@
$I_{c}(\rho)$ Closest incoherent state (43)
@@
Although the above-presented proofs employ several techniques that come from prior works [62, 63], the above proofs are not simple combinations of techniques that come from prior works. Since the two entries of the relative entropy and relative Rényi entropy are exchanged in our measure, the above proofs require new careful treatments for matrix convexity related to the parameter $\alpha$.
Remark 3
@@ VI Maximally correlated states
The discussion in Section II shows the operational meaning of $E_{\alpha}^{*}(\rho_{AB})$ including $E^{*}(\rho_{AB})$. We calculate $E^{*}(\rho_{AB})$ and $E^{*}_{\alpha}(\rho_{AB})$ for several special cases. When $\rho_{AB}$ is a pure state $|\phi\rangle\langle\phi|$ $E^{*}(\rho_{AB})$ and $E^{*}_{\alpha}(\rho_{AB})$ with $\alpha>1$ are infinity. However, $E^{*}_{\alpha}(\rho_{AB})$ with $\alpha\in[0,1)$ takes a finite value as follows. When the Schmidt coefficients of $|\phi\rangle\langle\phi|$ are $\sqrt{p_{i}}$ with $i=1,\ldots,d$, $\max_{\sigma_{AB}\in{\cal S}_{AB}}\langle\phi|\sigma|\phi\rangle=\max_{i}p_{i}$ [63, (40) and (97)]. Hence,
$e^{(\alpha-1)E_{\alpha}^{*}(\rho_{AB})}=\max_{\sigma_{AB}\in{\cal S}_{AB}}\operatorname{\mathrm{Tr}}\big{(}\rho^{\frac{1-\alpha}{2\alpha}}\sigma_{AB}\rho^{\frac{1-\alpha}{2\alpha}}\big{)}^{\alpha}$
$=$ $\max_{\sigma_{AB}\in{\cal S}_{AB}}\operatorname{\mathrm{Tr}}\big{(}|\phi\rangle\langle\phi|\sigma_{AB}|\phi\rangle\langle\phi|\big{)}^{\alpha}=\max_{\sigma_{AB}\in{\cal S}_{AB}}\langle\phi|\sigma_{AB}|\phi\rangle^{\alpha}$
$=$ $\Big{(}\max_{\sigma_{AB}\in{\cal S}_{AB}}\langle\phi|\sigma_{AB}|\phi\rangle\Big{)}^{\alpha}=(\max_{i}p_{i})^{\alpha}.$ (58)
@@
More generally, the calculation of these quantities can be converted to the calculation of the coherence measure for special states $\rho_{AB}$. Similar to the paper [63], we consider maximally correlated states. A state $\rho_{AB}$ on ${\mathcal{H}}_{A}\otimes{\mathcal{H}}_{B}$ is called a maximally correlated state when there exist bases $\{|u_{j,A}\rangle\}_{j}$ and $\{|u_{j,B}\rangle\}_{j}$ of ${\mathcal{H}}_{A}$ and ${\mathcal{H}}_{B}$ such that the support $\rho_{AB}$ is included in the subspace spanned by $\{|u_{j,A}\rangle|u_{j,B}\rangle\}_{j}$ on ${\mathcal{H}}_{A}\otimes{\mathcal{H}}_{B}$ [68, 69], [12, Section 8.8]. In this case, the state $\rho_{AB}$ is written as $\sum_{i,j}\theta_{i,j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{i,A}|\langle u_{i,B}|$. In this case, the calculation of $E^{*}(\rho_{AB})$ and $E^{*}_{\alpha}(\rho_{AB})$ can be simplified as follows. Similar to Theorem 1 in [63], the calculation of $E^{*}(\rho_{AB})$ and $E^{*}_{\alpha}(\rho_{AB})$ are reduced to coherence measures.
@@
where $\rho_{\theta}:=\sum_{i,j}\theta_{i,j}|j\rangle\langle i|$. The combination of Lemmas 9 and 10 guarantees that any maximally entangled state $\rho_{AB}$ satisfies
@@
$=\frac{1}{\alpha-1}\log\sum_{j=0}^{d-1}p_{j}^{1-\alpha}-\frac{\alpha}{\alpha-1}\log d$ (65)
for $\alpha\in(0,1)\cup(1,2]$. However, it is not easy to calculate these quantities for general states. Therefore, it is needed to develop an efficient method for their calculation in the general case.